Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2008-01-05
ISS On-Orbit Status 01/05/08

Saturday -- off-duty day for CDR Whitson, FE-1 Malenchenko and FE-2 Tani except for housekeeping and voluntary work.

CDR Whitson & FE-2 Tani started out with the daily reading of SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) experiment data accumulated during the night, for logging and filling in questionnaire entries in the SLEEP session file on the HRF-1 laptop for downlink. (To monitor the crewmembers' sleep/wake patterns and light exposure, Dan and Peggy wear a special Actiwatch device which measures the light levels encountered by him as well as his patterns of sleep and activity throughout the Expedition. The log entries are done within 15 minutes of final awakening for seven consecutive days, as part of the crew's discretionary 'job jar' task list.)

The crew performed the regular weekly three-hour task of thorough station cleaning. ("Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the Service Module (SM) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the FE's sleep station with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.)

As part of the house cleaning, FE-1 Yuri Malenchenko performed preventive maintenance cleaning on the FS5, FS6, VPkhO, VdPrK, VPrK & TsV2 fan grilles in the DC1 Docking Module and FGB (Funktsionalnyi-Grusovoi Blok).

For today's VolSci (Voluntary Weekend Science) program, Dan Tani completed an all-day session with the NASA/JSC experiment NUTRITION w/Repository, starting with the 24-hour urine collections at wakeup, to continue through tomorrow morning. Peggy Whitson also collected & processed blood samples from Dan. The samples were consecutively stored in the MELFI (Minus-Eighty Laboratory Freezer for ISS). (The current NUTRITION/Repository project is the most comprehensive in-flight study done by NASA to date of human physiologic changes during long-duration space flight. It includes measures of bone metabolism, oxidative damage, nutritional assessments, and hormonal changes, expanding the previous Clinical Nutritional Assessment profile (MR016L) testing in three ways: Addition of in-flight blood & urine collection (made possible by MELFI), normative markers of nutritional assessment, and a return session plus 30-day (R+30) session to allow evaluation of post-flight nutrition and implications for rehabilitation.)

For her VolSci part, CDR Whitson worked with the MSG (Microgravity Science Glovebox) and the InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions) experiment. (After activation of MSG and InSPACE & InSPACE-2 equipment, Peggy checked on alignment & focusing of the two MSG video cams, switched the magnetic field, changed out video recorder tapes and later deactivated InSPACE & MSG. InSPACE, conducted last in June 2006 by Jeff Williams on Increment 13, obtains basic data on magnetorheological fluids, i.e., a new class of "smart materials" that can be used to improve or develop new brake systems, seat suspensions robotics, clutches, airplane landing gear, and vibration damper systems. The dispersed particles are contained in CAs (Coil Assemblies) in the MSG that subject them to electric fields of certain strength and frequencies.)

The FE-1 meanwhile performed the routine servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. (Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers. Weekly SOZh reports (on Sundays) to TsUP/Moscow deal with number & dates of water and urine containers, counter readings of water consumption & urine collection, and total operating time of the POTOK air filtration system.)

The FE-2 performed his daily status check on the BCAT-3 (Binary Colloidal Alloy Test-3) science payload, running by itself in Node-2 since 12/13/07 (briefly interrupted for EVA-13 photo support). (The status check, conducted on the last image taken by the DCS 760 digital still camera which is controlled by EarthKAM software on an A31p laptop, is to verify proper image focus and camera alignment. The SSC (Station Support Computer) is taking photography of the phase separation occurring in the BCAT Sample 3, with the photo flash going off every half hour. Special uplink to Dan this morning: 'The BCAT team downloaded the new images and reports that they look fabulous. The team is very excited and looking forward to pressing ahead and making more progress in coming days.')

With the Russian O2 (oxygen) generator 'Elektron' turned off since 12/28/07, Yuri had time reserved for another 1-hour O2 repress of the cabin atmosphere from Progress M-62/27P storage tankage, if required. Yesterday's repress added about 8 mmHg of O2 into the ISS as per plan. (The Elektron will remain powered down until 1/9/08 to conserve hardware lifetime. During this time, the station will be periodically repressurized with oxygen from Progress 27P.)

The crewmembers performed their regular 2.5-hr physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (CDR), TVIS treadmill (FE-1, FE-2), RED (CDR, FE-2) and VELO bike with bungee cord load trainer (FE-1).

Afterwards, Peggy copied the exercise data file to the MEC (Medical Equipment Computer) laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

Working off his voluntary 'time available' task list, Yuri took photographs of the externally installed SKK-9/SM and SKK-2/DC1 replaceable materials exposure experiments. (SKK-9 was to be photographed through the DC1 EVA hatch window, SKK-2 from the SM window #6.)

A second job item on the FE-1's discretionary list for today was another KPT-3 session to make observations and take aerial KPT-3 photography of environmental conditions for Russia's Environmental Safety Agency (ECON) using the Nikon D2X digital camera with SIGMA 300-800mm telephoto lens. (Targets today were contamination areas in the Seattle area and in the Pacific Ocean.)

At ~7:45am EST, Yuri participated in an interview by journalist Ekaterina Beloglazova, a correspondent of Rossiyskiy Kosmos Magazine, responding to questions uplinked beforehand. ('We know that because of the Shuttle launch delay your program underwent some changes. What kind of experiments, operations and observations did you perform? What are your impressions and do you have some interesting results?'; 'Tell us about the new European module Columbus.'; 'How much is this expedition different from your past missions?'; 'Tell us about your colleagues, your work, your relations, your habits. What is your daily schedule, where do you live, what do you like, how do you spend your free time, etc.?')

At ~9:10am EST, the crewmembers held their regular weekly planning conference (WPC) with the ground, discussing next week's "Look-Ahead Plan" (prepared jointly by MCC-H and TsUP/Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

At ~9:40am, the FE-1 had his weekly PFC (Private Family Conference) via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on the SSC-9 laptop).
The crew worked out in their regular 2.5-hr. physical exercise program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (CDR, FE-2), TVIS treadmill (FE-1), RED resistive exerciser (CDR, FE-2), and VELO bike with bungee cord load trainer (FE-1).

SKV2 Update: After the recent successful SKV2 air conditioner reactivation, Malenchenko on 1/3 found condensate water behind some SM panels (405, 407, 409, etc). Cause of the overflow was determined to be a faulty 'full' indicator pressure sensor inside the SBK condensate collection tank which failed to shut down the NOK2 condensate pump. After SKV2 deactivation by TsUP, Yuri replaced the tank with a new spare, reactivated the air conditioner and mopped up the water spills. (SBK has 9 l water capacity and contains an absorptive material inside the tank. There are 6 spares remaining on board. These SBK tanks typically last 1 to 2 years.)

Weekly Science Update (Expedition Sixteen -- Week 11)

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): FE-1 performed the exchange of memory card on 12/27. Radiation measurements continue nominally in the PIRS module. Next activities (rotation of the ALTEINO instrument, replacement of memory card and downlink) are currently scheduled on 1/8 and 1/9.

ANITA: Completed.

BCAT (Binary Colloidal Alloy Test): Reserve.

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): 'Dan, we are waiting to get the new launch date so we can schedule your second in-flight session to meet the R-21 to R-14 requirement. This activity will be hard scheduled when we get a firm launch date.'

CFE (Capillary Flow Experiment): Reserve.

CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): In progress.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): In progress.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

ELITE-S2 (Elaboratore Immagini Televisive - Space 2): ELITE IMU check-out was successful. The ground received good health & status data and could send commands to the payload. The payload reply was nominal and at the end the payload was shut down via ground command.

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): In progress.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): 'InSPACE-2 setup was completed on 1/4 and we were correctly configured for today's (1/5) first run. The crew provided us with different video views and close-ups of the hardware setup to help us assess the hardware on the ground.'

Integrated Immune: In progress.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): 'Thanks for operating LOCAD-PTS during 12/21-24. The LOCAD team is extremely grateful for your hard work in performing all the scheduled swabs (and more!) and excited about the results. To give you some feedback, here's a brief summary. Dan performed a valuable survey of 12 different sites on ISS during his two sessions. The sites he chose were excellent, and just what we were looking for based on the criteria sent up before the sessions. Some of these sites gave much higher endotoxin readings (e.g. OFFSU bathroom mirror 3.07 endotoxin units (EU)/ml, SM table top 0.564 EU/ml, and CEVIS handle 0.469), than we had seen previously (maximum reading of 0.445 EU/ml out of 23 swabs performed by Suni). Peggy's session helped further our operational database for comparison of LOCAD-PTS data with that of the CHeCS SSK culture methods. From the data we have so far, the sites that Peggy analyzed look clean (i.e. low EU/ml) and it will be interesting to see what SSK culture we get from those same sites. OPERATIONAL ANOMALIES: 1. Frozen display/'temp too high' reading. We had mistakenly informed Dan that he could leave the cartridge in the reader indefinitely before returning to retrieve the data. In actual fact, it appears the inserted cartridge caused the internal temperature of the reader to rise above the optimal 37 Celsius, to 40-43 Celsius. SOLUTION: We will note in procedures to leave the cartridge in the reader no longer than 30 minutes. 2. Air bubbles in water cartridge and during mixing/dispensing. You both worked very well through procedures to minimize air bubbles and accurately dispense 4 droplets of equal volume in the LOCAD-PTS cartridge. However, we acknowledge that air bubbles are occurring - even before use- and that they complicate dispensing. The cause of air bubbles in the water cartridges may be partly due to inadequately sealed swabbing kit bags. SOLUTIONS: Swabbing kit bags scheduled for flight 1J/A have been more thoroughly heat-sealed. We believe this will reduce the rate of evaporation and bubble formation in the water cartridge. Introduction of bubbles during mixing is difficult to avoid, but we will revisit procedures to see if we can minimize bubbles still further.'

MISSE (Materials ISS Experiment): Ongoing.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 'Pirs'.

MULTIGEN-1: MULTIGEN-1 samples will be downloaded on STS-122 (1E).

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-2 (Nitric Oxide Analyzer): Planned.

NUTRITION/REPOSITORY: 'Peggy did a great job completing the FD60 Nutrition/Repository run. Blood operations were completed in record time'.

PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): 'Peggy, we were able to downlink the data from your last download/initialization session and send it to the PI. She states that your data continues to look great. Thanks for the additional Sleep logging. You continue to go above and beyond the requirements and the PI greatly appreciates it. Your remaining activities off the task list include: monthly downloads, two more weeks of target sleep logging with sleep shifting, and doffing the Actiwatch.' -- 'Dan, we were able to downlink the data from your last download/initialization session and send it to the PI. She states that your data continues to look great. Please continue to Sleep log this week to meet the three week requirement. We have left this activity on your task list per your suggestion. Your remaining activities off the task list include: at least one more download/initialization session (depending on shuttle launch), one more week of sleep logging, and doffing the Actiwatch on the Shuttle.'

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): In progress.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.

CEO (Crew Earth Observation): Through 1/2 the ground has received a total of 8,323 frames of ISS/CEO imagery for review and cataloging. 'As our staff returns from holidays and vacations we note with delight that you continued to be an active and enthusiastic Earth observer. Over three thousand frames were acquired since mid-December including frames with times corresponding to 48 of our target requests. We will begin providing feedback on this large backlog in the coming weeks. We will also look at the first 225 frames with D1X and D2Xs cameras and give you an assessment. Your colorful, landmark photo of the São Simão Reservoir in east-central Brazil represents the 300,000th image of the Earth downlinked from the ISS. Congratulations! It will be published on NASA/GSFC's Earth Observatory website this weekend.'

CEO photo targets uplinked for today again were Ganges River Delta, Asia (weather was predicted to be clear over the Ganges delta area. A nadir mapping pass along-track was requested to capture the current configuration of outlet channels and vegetation patterns), Irrawaddy River Delta, Burma (clear weather was predicted to continue over the Irrawaddy River delta. A nadir mapping pass along track was requested to track changes to the outlet channels, island morphology, and vegetation patterns), Polar Mesospheric Clouds - PMC, Antarctica (IPY--PMC radar research station active. GMTs for this and PMC opportunities uplinked have been chosen for closeness to the Antarctic PMC radar research site at 73S 13W. Radar is switched on during ISS passes at the GMTs given. But the crew was to feel free to look south during any night awake pass), and Gulf of Fonseca, Central America (ISS had a nadir pass over this Central America embayment. High resolution mapping along the coastline was requested to document land cover and land use change).

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use